OpenStudy (anonymous):

what is the geometric mean of 7 and 12

OpenStudy (idkwut):

Love.

OpenStudy (idkwut):

Lulz, jk. b^2=a*c b^2=7*12 b^2=84 b=84^1/2 = ?? Try it.

OpenStudy (jmark):

Geometric mean of two term a,b is given by √ab mean of 7 and 12 is =√7x12 = √84 ICSE Results 2014 - http://goo.gl/aZye9y